The Mathematics Club at Einstein High School noticed that the school's
student of the week was chosen using an arithmetic sequence. The name
of every student in the school was listed alphabetically and numbered.
The numbers of the first 5 students chosen are listed below.
21, 26, 31, 36, 41
What will be the number of the 111th student chosen?

Answer by stanbon(75887)

The name
of every student in the school was listed alphabetically and numbered.
The numbers of the first 5 students chosen are listed below.
21, 26, 31, 36, 41
What will be the number of the 111th student chosen?
----
a(1) = 21
d = 26-21 = 5
------
a(111) = 21 + 110*5 = 21 + 550 = 571
-------------
Cheers,
Stan H.
---------------